# HyperGaussians: High-Dimensional Gaussian Splatting for High-Fidelity Animatable Face Avatars

### v2

@@ Abstract
We introduce HyperGaussians, a novel extension of 3D Gaussian Splatting for high-quality animatable face avatars. Creating such detailed face avatars from videos is a challenging problem and has numerous applications in augmented and virtual reality. While tremendous successes have been achieved for static faces, animatable avatars from monocular videos still fall in the uncanny valley. The de facto standard, 3D Gaussian Splatting (3DGS), represents a face through a collection of 3D Gaussian primitives. 3DGS excels at rendering static faces, but the state-of-the-art still struggles with nonlinear deformations, complex lighting effects, and fine details. While most related works focus on predicting better Gaussian parameters from expression codes, we rethink the 3D Gaussian representation itself and how to make it more expressive. Our insights lead to a novel extension of 3D Gaussians to high-dimensional multivariate Gaussians, dubbed ’HyperGaussians’. The higher dimensionality increases expressivity through conditioning on a learnable local embedding. However, splatting HyperGaussians is computationally expensive because it requires inverting a high-dimensional covariance matrix. We solve this by reparameterizing the covariance matrix, dubbed the ’inverse covariance trick’. This trick boosts the efficiency so that HyperGaussians can be seamlessly integrated into existing models. To demonstrate this, we plug in HyperGaussians into the state-of-the-art in fast monocular face avatars: FlashAvatar. Our evaluation on 19 subjects from 5 face datasets shows that HyperGaussians outperform 3DGS numerically and visually, particularly for high-frequency details like eyeglass frames, teeth, complex facial movements, and specular reflections.
@@
Figure 1: We propose a novel representation for monocular face avatars: HyperGaussians. HyperGaussians extend 3D Gaussians to higher dimensions, resulting in improved high-frequency details for specular reflections and thin structures. HyperGaussians can be plugged into existing models with minimal overhead. This figure shows the effect of plugging HyperGaussian (bottom) into the state-of-the-art for monocular face avatars, FlashAvatar [51] (top). Note the improvement in specular reflections, complex deformations, and thin structures.
@@
Figure 2: We propose an extension to 3D Gaussians, dubbed HyperGaussians, and plug them into an existing method for face avatars, FlashAvatar [51]. FlashAvatar modulates 3D Gaussian primitives with expression-dependent offsets $\Delta$. We make a single modification to the pipeline: We plug in HyperGaussians (Sec. 3.2) between the MLP output and the rasterization, which modifies the offsets $\Delta$ in higher dimensions. Instead of directly predicting offsets $\Delta$, we predict a latent $z_{\psi}$ that conditions HyperGaussians. Without any other modifications or hyperparameter tuning, this simple change enhances details in the final avatar (Fig. 7) and leads to a performance boost (Tab. 2). This figure has been adapted from FlashAvatar [51].
@@ 1 Introduction
Modelling dynamic scenes like human faces is a long-standing problem in 3D Computer Vision and has a variety of applications in augmented and virtual reality, the movie and gaming industry, and virtual telepresence [20, 36]. Faces exhibit highly nonlinear deformations with topology changes, such as eye blinks and mouth opening. In addition, faces show complex lighting effects like specular reflections in the eyes and on glasses [23, 24, 43, 45]. It is important to solve these details and cross the uncanny valley, enhancing human-avatar interaction, and creating truly lifelike avatars [35, 17]. One crucial element towards this goal is the development of high-quality and efficient 3D representations.
@@
Recently, 3D Gaussian Splatting [18] has become the de facto standard for modeling humans due to its quality and efficiency [46, 8, 51, 63, 43]. The state-of-the-art for human faces rigs Gaussians with a Morphable Face Models [2, 13, 27]. Each Gaussian is attached to a mesh triangle and follows its deformation, for example, using linear blend-skinning [27, 29, 22]. Such linear models are a good approximation, but they cannot represent nonlinear deformations and specular effects. To mitigate this, recent works leverage Neural Networks to predict offsets based on an input expression [51, 46, 8]. This improves results over static Gaussians, but it is not sufficient to model details. Fig. 7 shows how these methods struggle to represent thin structures like teeth, hair, and glass frames (rows 1 - 4) and specular reflections on eyes and glasses (rows 1 and 3), and nonlinear deformations like closing eyelids (rows 2 and 4). These high-frequency details are key for crossing the uncanny valley.
@@
In this work, we extend 3D Gaussian Splatting [18] to arbitrary higher dimensions and call the novel representation HyperGaussians (Sec. 3.2). HyperGaussians are an expressive, lightweight, and fast high-dimensional representation for high-quality face avatars. At their core, they are multivariate Gaussians. The additional dimensions and the local embeddings improve complex local deformations and specular effects. However, due to the higher dimensionality, a naïve implementation of high-dimensional Gaussians requires extremely high computational resources for splatting (Fig. 5). To solve this, we propose a re-parameterization, called the inverse covariance trick (Sec. 3.2). With the inverse covariance trick, high-dimensional Gaussian primitives can be splatted efficiently in real-time.
@@
To the best of our knowledge, we are the first to improve high-frequency details in facial avatars by enhancing the low-level parameterization of the Gaussian primitives. Most related works focus on improving a model architecture or training scheme [51, 8, 46, 42, 21] that estimates the dynamics of the Gaussian parameters using neural networks. Our contribution, HyperGaussians, is orthogonal to these efforts. HyperGaussian builds on top of the 3DGS framework and thus can be readily plugged into an existing pipeline and boost its performance out-of-the-box, without any other changes to the model, architecture, or hyperparameters. We demonstrate this by integrating our proposed HyperGaussians into the state-of-the-art in fast face avatar learning from monocular videos, FlashAvatar [51]. The only change we make is to replace the 3D Gaussians with HyperGaussians, the rest of the model remains exactly the same. We evaluate on videos of $19$ subjects from $4$ different datasets. The result is a boost in performance for self- and cross-reenactment, as demonstrated in Figs. 1, 7, 8 and 3 and Tab. 2. HyperGaussian can model thin structures like glass frames and teeth, specular reflections on eyes and glasses, and complex nonlinear deformations such as closing of the eyelid.
@@
  • A novel expressive representation for modeling high-frequency and dynamic effects, dubbed HyperGaussians.
  • The inverse covariance trick, a technical contribution that dramatically increases the computational efficiency of high-dimensional Gaussian conditioning.
  • A case study on $19$ subjects from $4$ datatasets, where we plug in HyperGaussian into the state-of-the-art fast face avatar, FlashAvatar, demonstrating improved rendering quality without any other changes to the model.
FlashAvatar
Ours
5 seconds 20 seconds 60 seconds 180 seconds 600 seconds
Figure 3: We plug in our proposed HyperGaussian into FlashAvatar [51] and compare the convergence speed. The only difference between FlashAvatar and Ours is the substitution of 3D Gaussians (top) with HyperGaussians (bottom), as described in Sec. 3.3. HyperGaussians displays sharper results throughout the training.
@@ 2 Related Work
Traditional 3D representations, such as meshes and point clouds, are increasingly being replaced by more innovative methods. Neural Radiance Fields (NeRFs) [33], Neural Implicit Functions [37, 32], and 3D Gaussians [18] can represent 3D scenes with a high level of detail and enable photorealistic novel view synthesis. Neural Radiance Fields and Implicit Functions can be adapted to capture dynamic effects by conditioning on a time dimension [38, 39, 41] to enable the replay of videos from novel views and featuring moving objects or a change in facial expressions. NeRFies [38] optimizes a continuous deformation field that maps from a deformed to a canonical space but suffers from artifacts when topological changes arise, e.g., a mouth opening. HyperNeRF [39] showed that these issues can be mitigated by modeling the deformation field in higher dimensions. The HyperNeRF deformation field slices the high-dimensional deformation field with learnable embeddings. Inspired by HyperNeRF, we propose an extension to 3D Gaussian Splatting by modeling Gaussians in a higher-dimensional space.
N-dimensional Gaussian Splatting
A number of works have explored Gaussian primitives with non-3D dimensionalities [16, 50, 10, 15, 60, 30, 56]. 1D Cylindrical Gaussian Splatting is being used for hair modeling [60, 30, 56]. 2D Gaussian Splatting [16] uses oriented planar Gaussian disks to improve geometric consistency in Gaussian Splatting. 4D Gaussian Splatting [50] extends 3DGS for rendering dynamic scenes. They propose a custom CUDA kernel for splatting 4D Gaussians with a time dimension. This approach reconstructs detailed videos, but it does not support higher dimensions. NDGS [10] formulates a Gaussian Mixture Model for representing static scenes with high appearance variability, which is effective for highly reflective surfaces. Their representation exhibits some key differences from ours. Their representation has no degrees of freedom for the conditional orientation of the 3D Gaussians, i.e., the Gaussians cannot rotate. Moreover, the rendered size of a Gaussian depends on the probability density function of the joint distribution. This causes Gaussians to disappear under large deformations and leads to semantic subparts of the scene being modeled by multiple Gaussians, which are invisible most of the time. Our formulation removes these limitations and ensures that the Gaussians deform consistently.
@@
A popular goal is reconstructing a face avatar with a high level of fidelity and rendering it under a novel pose and expression [25, 6, 31, 9, 26, 3, 5, 4, 58]. 3D Gaussians have become the de facto standard for modeling such face avatars [52, 25, 44, 31, 49, 54, 48, 59]. Tab. 1 provides an overview of the most closely related works. The state-of-the-art [51, 46, 21, 42] attaches Gaussians to the mesh of a Morphable Model (3DMM) [13, 2, 27]. 3DMMs serve as a strong shape prior and were already extensively made use of in implicit avatars, such as NerFACE [11] and INSTA [61]. This enables driving the avatar with controlled expressions and head poses with linear blend skinning [22, 27]. SplattingAvatar [46] improves over vanilla blendskinning by optimizing embeddings on the mesh. However, SplattingAvatar does not consider that deformations also affect the appearance, for example, a head pose change will move the specular reflections on glasses (Fig. 7), leading to a blurry rendering. Recent works take a step towards improving such high-frequency effects by predicting expression-dependent Gaussian parameter offsets [51, 63, 9, 8], allowing the Gaussian parameters to change for different expressions. Predicting these expression-dependent offsets makes FlashAvatar [51] the current state-of-the-art for fast face avatar modeling.
Method Representation Local context Dynamic context
GaussianAvatars [42] 3DGS $\times$ $\times$
SurFHead [21] 2DGS $\times$ $\times$
SplattingAvatar [46] 3DGS $\checkmark$ $\times$
MonoGaussianAvatar [8] 3DGS $\times$ $\checkmark$
FlashAvatar [51] 3DGS $\times$ $\checkmark$
FlashAvatar w/ HyperGaussian (Ours) HyperGaussian $\checkmark$ $\checkmark$
Table 1: We show the differences to the most closely related works. GaussianAvatars [42] and SuRFHead [21] deform 3D Gaussians based on an underlying FLAME mesh [27] without local embeddings or dynamic inputs like facial expressions. SplattingAvatar [46] optimizes local embeddings but the Gaussian properties are not dependent on expressions or pose. MonoGaussianAvatar [8] and FlashAvatar [51] predict expression-dependent offsets to the Gaussian properties, but their lack of local context leads to blurry or distorted results, see Fig. 7. Our proposed representation attaches high-dimensional Gaussians to the mesh and optimizes learnable local embeddings for modulating the Gaussian properties based on expressions.
@@ 3 Method
We propose a novel representation for modeling dynamic 3D scenes and apply it to face avatars. Our novel representation extends 3D Gaussians [18] to higher dimensions and extends each Gaussian primitive with an embedding that provides local context. This enables representing finer details such as a thin frame of glasses, tiny gaps between teeth, specular reflections on the eyes, and glasses (Fig. 1). We dub our novel representation HyperGaussians.
We start with preliminaries in Sec. 3.1 and introduce the general framework of HyperGaussians in Sec. 3.2. While already performing well for a low-dimensional hyperspace (less than 5 hyper-space dimensions, Tab. 3), a naïve implementation does not scale because of higher computational and memory load (Fig. 5). As a solution, we propose the inverse covariance trick. We demonstrate the effectiveness of HyperGaussians to reconstruct detailed face avatars from monocular videos in Sec. 3.3. Specifically, we plug in HyperGaussians into an existing method, FlashAvatar [51]. A simple replacement of 3D Gaussians with HyperGaussians without any hyperparameter-tuning boosts FlashAvatar’s performance (Tab. 2).
@@ 3.1 Preliminary: 3D Gaussian Splatting
3D Gaussian Splatting [18] (3DGS) models a static scene with colored anisotropic 3D Gaussians. The Gaussians are parameterized by their mean ${\boldsymbol{\mu}\in\mathbb{R}^{3}}$ and covariance matrix ${\boldsymbol{\Sigma}\in\mathbb{R}^{3\times 3}}$ and their appearance is represented by their opacity $\alpha$ and a color $\boldsymbol{c}$, where common choices are RGB color or Spherical Harmonics for view-dependent effects. Given images, camera parameters, and a sparse estimated point cloud, 3DGS can reconstruct a scene by optimizing a set of Gaussians via differentiable rasterization.
To ensure that the covariance matrices remain positive semi-definite during optimization, 3DGS [18] first defines a parametric ellipsoid using a scaling matrix $\boldsymbol{S}$ and a rotation matrix $\boldsymbol{R}$, then constructs the covariance matrix as
$\boldsymbol{\Sigma}=\boldsymbol{R}\boldsymbol{S}\boldsymbol{S}^{\top}\boldsymbol{R}^{\top}.$ (1)
These matrices are themselves parameterized by a scaling vector ${\boldsymbol{s}\in\mathbb{R}^{3}}$ and a unit quaternion ${\boldsymbol{q}\in\mathbb{R}^{4}}$, respectively. We denote the set of Gaussians as $\{(\boldsymbol{\mu}_{i},\boldsymbol{s}_{i},\boldsymbol{q}_{i},\boldsymbol{c}_{i},\alpha_{i})\}_{i=1}^{n}$.
Novel views can be rendered by splatting the Gaussians, following these two steps: the Gaussians are first projected onto the camera plane and then alpha blended together with respect to their evaluated density $G(\boldsymbol{x})=\exp\bigl{(}-\frac{1}{2}(\boldsymbol{x}-\boldsymbol{\mu})^{\top}\boldsymbol{\Sigma}^{-1}(\boldsymbol{x}-\boldsymbol{\mu})\bigr{)}$ and learned opacity $\alpha$. For each pixel, its color can then be computed using alpha blending:
$\boldsymbol{C}=\sum_{i}\boldsymbol{c}_{i}^{\prime}\alpha_{i}^{\prime}\prod_{j=1}^{i-1}(1-\alpha_{j}^{\prime}).$ (2)

### v3

@@ Abstract
We introduce HyperGaussians, a novel extension of 3D Gaussian Splatting for high-quality animatable face avatars. While tremendous successes have been achieved for static faces, animatable avatars from dynamic videos still fall in the uncanny valley. The de facto standard, 3D Gaussian Splatting (3DGS), represents a face through a collection of 3D Gaussian primitives. 3DGS excels at rendering static faces, but the state-of-the-art still struggles with nonlinear deformations, complex lighting effects, and fine details. While most related works focus on predicting better Gaussian parameters from expression codes, we rethink the 3D Gaussian representation itself and how to make it more expressive. Our insights lead to a novel extension of 3D Gaussians to high-dimensional multivariate Gaussians, dubbed ’HyperGaussians’. The higher dimensionality increases expressivity through conditioning on a learnable local embedding. However, splatting HyperGaussians is computationally expensive because it requires inverting a high-dimensional covariance matrix. We solve this by reparameterizing the covariance matrix, dubbed the ’inverse covariance trick’. This trick boosts the efficiency so that HyperGaussians can be seamlessly integrated into existing models. To demonstrate this, we plug in HyperGaussians into two state-of-the-art methods for face avatars: FlashAvatar and GaussianHeadAvatar. Our evaluation on 29 subjects from 6 face datasets shows that HyperGaussians outperform 3DGS numerically and visually, particularly for high-frequency details like eyes, teeth, wrinkles, and specular reflections.
@@
Figure 1: We propose a plug-and-play enhancement for 3D Gaussians: HyperGaussians. HyperGaussians extend 3D Gaussian Splatting to higher dimensions, resulting in improved high-frequency details for specular reflections and thin structures. HyperGaussians can be easily integrated into existing models with minimal overhead. This figure shows the effect of plugging HyperGaussians (bottom) into the state-of-the-art for multi-view face avatars, GaussianHeadAvatar [71] (top). Note the improvement in specular reflections, complex deformations, and fine details. Fig. 4 shows results for monocular face avatars [70].
@@
Figure 2: We propose an expressive extension to 3D Gaussians, dubbed HyperGaussians, and plug them into existing methods for face avatars, FlashAvatar [70] and GaussianHeadAvatar [71]. FlashAvatar modulates 3D Gaussian primitives with expression-dependent offsets $\Delta$. We make a single modification to the pipeline: plugging HyperGaussians (Sec. 3.2) in between the MLP output and the rasterization, which modifies the offsets $\Delta$ in higher dimensions. Instead of directly predicting offsets $\Delta$, we predict a latent $z_{\psi}$ that conditions HyperGaussians. Without any other modifications or hyperparameter tuning, this simple change leads to a performance boost in rendering high-frequency details in the final avatar (Tab. 1 and Fig. 4). This figure has been adapted from FlashAvatar [70].
@@ 1 Introduction
Modeling dynamic scenes, such as human faces, is a long-standing problem in 3D Computer Vision and has a variety of applications in augmented and virtual reality, the entertainment industry, and virtual telepresence [30, 50]. Faces exhibit highly nonlinear deformations such as eye blinks and mouth opening, and complex lighting effects like specular reflections in the eyes and on glasses [33, 34, 58, 59]. Faithfully modeling these details is essential to overcome the uncanny valley, enhancing human-avatar interaction, and creating truly lifelike avatars [48, 25]. One crucial element towards this goal is the development of high-quality and efficient 3D representations.
@@
Recently, 3D Gaussian Splatting [26] has emerged as the standard for modeling humans due to its quality and efficiency [60, 10, 70, 81, 58, 71]. The state-of-the-art for human faces, rigs Gaussians with Morphable Face Models [3, 20, 38]. Each Gaussian is attached to a mesh triangle and follows its deformation, for example, using linear blend skinning [38, 42, 32]. Such linear models are a good approximation, but they cannot represent nonlinear deformations and specular effects. To mitigate this, recent work leverages neural networks to predict offsets from an input expression [70, 60, 10, 71], improving results over static Gaussians. Yet, Fig. 4 and Fig. 6 show that these methods struggle to represent thin structures like teeth and hair, specular reflections on eyes and glasses, and nonlinear deformations such as closing eyelids. These high-frequency details are key to achieving believable digital human avatars.
@@
In this work, we extend 3D Gaussian Splatting [26] to arbitrary higher dimensions and call the novel representation HyperGaussians (Sec. 3.2). HyperGaussians are an expressive, high-dimensional representation for high-quality face avatars. We motivate HyperGaussians by analyzing them through a probabilistic lens (Sec. 3.2). At their core, they are multivariate Gaussians that parameterize a conditional (with respect to local embeddings) distribution over vanilla 3DGS scenes. The additional dimensions and local embeddings improve complex local deformations and specular effects. However, due to the higher dimensionality, a naïve implementation of high-dimensional Gaussians requires substantial computational resources for splatting (Fig. 3). To solve this, we propose a re-parameterization, called the inverse covariance trick (Sec. 3.2). Using the inverse covariance trick, high-dimensional Gaussian primitives can be splatted efficiently in real time at 300 FPS.
@@
To the best of our knowledge, we are the first to improve high-frequency details in facial avatars by enhancing the low-level parameterization of the Gaussian primitives. Most related works focus on improving a model architecture or training scheme [70, 10, 60, 57, 31, 71] that estimates the dynamics of the Gaussian parameters using neural networks. Our contribution, HyperGaussians, is orthogonal to these efforts. HyperGaussians build on top of the 3DGS framework and thus can be readily plugged into an existing pipeline and boost its performance out-of-the-box, without any other changes to the model, architecture, or hyperparameters. We demonstrate this by integrating our proposed HyperGaussians into the state-of-the-art in fast face avatar learning from monocular video, FlashAvatar [70], and the state-of-the-art for multi-view video, GaussianHeadAvatar [71]. The only change required is to replace the 3D Gaussians with HyperGaussians, the rest stays exactly the same. We evaluate on videos of $29$ subjects from $6$ different datasets. The result is a boost in performance for self- and cross-reenactment, as demonstrated in Figs. 1, 4, 5 and 6 and Tab. 1. Specifically, HyperGaussians significantly improve thin structures like glass frames and teeth, specular reflections on eyes and glasses, and complex nonlinear deformations such as closing of the eyelid and wrinkles.
@@
  • HyperGaussians, a novel, expressive representation for modeling high-frequency and dynamic effects supported by a probabilistic interpretation.
  • The inverse covariance trick, a technical contribution that yields a substantial improvement in computational efficiency of high-dimensional Gaussian conditioning.
  • Evaluation of HyperGaussians applied to the state-of-the-art for fast face avatars, FlashAvatar, and multi-view avatars, GaussianHeadAvatar, demonstrating improved rendering quality without any other changes across $29$ subjects from $6$ datasets.
@@ 2 Related Work
Neural 3D representations like Neural Radiance Fields (NeRFs) [46], Neural Implicit Functions [51, 45], and 3D Gaussians [26] can represent 3D scenes with a high level of detail and enable photorealistic novel view synthesis. NeRFs and Implicit Functions can be conditioned to model dynamic effects over time [52, 53, 56] or facial expressions [15]. NeRFies [52] optimizes a continuous deformation field that maps from a deformed to a canonical space, but it suffers from artifacts when topological changes occur, e.g., a mouth opening. HyperNeRF [53] shows that these issues can be mitigated by modeling the deformation field in higher dimensions and slicing through it using learnable embeddings. Inspired by HyperNeRF, we propose an extension to 3D Gaussian Splatting by modeling Gaussians in a higher-dimensional space.
Gaussian Splatting
3D Gaussian Splatting (3DGS) [26] has effectively become the standard for modeling static scenes. Several works have explored extensions to non-standard dimensionalities [24, 69, 13, 22, 78, 43, 74]. 1D Cylindrical Gaussian Splatting is being used for hair modeling [78, 43, 74]. 2DGS [24] uses oriented planar Gaussian disks to improve geometric consistency in Gaussian Splatting. 4DGS [69] extends 3DGS for rendering dynamic scenes by proposing a custom CUDA kernel for splatting 4D Gaussians with a time dimension. NDGS [13] formulates a Gaussian Mixture Model for representing static scenes with high appearance variability, which is effective for highly reflective surfaces. Particular instances of NDGS, such as 6DGS [18] and 7DGS [19], use the additional dimensions to augment the Gaussians with view direction and time. We propose a novel, unified probabilistic interpretation of these efforts and identify critical limitations in their representation (Sec. 3.2). We address these shortcomings and highlight two key differences between their representations and HyperGaussians. First, their representation has no degrees of freedom for the conditional orientation of the 3D Gaussians, i.e., the Gaussians cannot rotate. Second, the rendered size of a Gaussian depends on the probability density function of the joint distribution. This causes Gaussians to disappear under large deformations, leading to semantic subparts of the scene being modeled by multiple Gaussians that are invisible most of the time. Our formulation removes these limitations and ensures that the Gaussians deform consistently. Furthermore, NDGS-based methods are limited by the computational effort required to condition their primitives. This makes them unsuitable for high-dimensional applications. Our inverse covariance trick overcomes this issue, as shown in Fig. 3, and thus provides a more scalable framework. We compare with NDGS in the supplementary material.
@@
A popular goal is reconstructing a face avatar [36, 9, 44, 12, 35, 6, 4, 5, 76, 29, 1] with high fidelity and rendering it under a novel pose and expression. Most recent methods [70, 71, 36, 58, 44, 66, 72, 63, 77, 1, 37, 14, 68] render avatars with Gaussian Splatting. The state-of-the-art [70, 60, 31, 57, 71, 1, 37, 14, 68] attaches Gaussians to the mesh of a Morphable Model (3DMM) [20, 3, 38]. 3DMMs serve as strong shape priors and have already been extensively used in implicit avatars, such as NerFACE [15] and INSTA [79]. This enables driving the avatar with controlled expressions and head poses through linear blend skinning [32, 38]. SplattingAvatar [60] improves over vanilla blend skinning by optimizing embeddings on the mesh. However, SplattingAvatar does not account for the fact that deformations also affect appearance. For example, a change in head pose will displace the specular reflections on glasses (Fig. 4), leading to a blurry rendering. This can be improved by predicting expression-dependent Gaussian parameter offsets [70, 81, 12, 10, 71, 1], allowing the Gaussian parameters to change for different expressions.
In highly sparse settings, recent work proposes large, pre-trained priors for predicting Gaussian attributes in a feed-forward manner or through pseudo ground-truth fitting using multi-view generative models [23, 11, 39, 64, 29, 65, 7]. However, the resource-intensive pre-training renders adaptations of these methods infeasible in low-compute environments. We therefore focus on optimization-based techniques in our analysis.
In this paper, we boost high-frequency details in state-of-the-art face avatars [70, 71] by replacing the vanilla 3D Gaussians [26] with a higher-dimensional representation.
@@ 3 Method
We propose a novel representation for modeling dynamic 3D scenes and apply it to face avatars. Our novel representation extends 3D Gaussians [26] to higher dimensions and augments each Gaussian primitive with the ability to adapt according to local embeddings. This enables the representation of finer details, such as specular reflections on glasses, skin wrinkles, and text labels (Fig. 1). We dub our novel representation HyperGaussians.
In Sec. 3.1, we reiterate on 3DGS before introducing HyperGaussians in Sec. 3.2, where we additionally present the Bayesian view, as well as the inverse covariance trick to massively reduce compute and memory load, making HyperGaussians highly scalable. Sec. 3.3 shows how we integrate HyperGaussians into the existing methods FlashAvatar [70] and GaussianHeadAvatar [71] to boost their performance (Tab. 1) with minimal modifications.
@@ 3.1 Preliminary: 3D Gaussian Splatting
3D Gaussian Splatting (3DGS) [26] models a static scene with colored anisotropic 3D Gaussians. The Gaussians are parameterized by their mean $\boldsymbol{\mu}_{\mathrm{3D}}\in\mathbb{R}^{3}$} and covariance matrix $\boldsymbol{\Sigma}_{\mathrm{3D}}\in\mathbb{R}^{3\times 3}$} and their appearance is represented by their opacity $\alpha$ and a color $\boldsymbol{c}$, where common choices are RGB color or Spherical Harmonics for view-dependent effects. Given images, camera parameters, and a sparse point cloud estimate, 3DGS can reconstruct a scene by optimizing a set of Gaussians via differentiable rasterization.
To ensure that the covariance matrices remain positive semi-definite during optimization, 3DGS first defines a parametric ellipsoid using a scaling matrix $\boldsymbol{S}$ and a rotation matrix $\boldsymbol{R}$, then constructs the covariance matrix as
$\boldsymbol{\Sigma}_{\mathrm{3D}}=\boldsymbol{R}\boldsymbol{S}\boldsymbol{S}^{\top}\boldsymbol{R}^{\top}.$ (1)
These matrices are themselves parameterized by a scaling vector $\boldsymbol{s}\in\mathbb{R}^{3}$} and a unit quaternion $\boldsymbol{q}\in\mathbb{R}^{4}$}, respectively. We denote the set of Gaussians as $\{(\boldsymbol{\mu}_{\mathrm{3D}}^{i},\boldsymbol{s}^{i},\boldsymbol{q}^{i},\boldsymbol{c}^{i},\alpha^{i})\}_{i=1}^{n}$.
Novel views can be rendered by splatting the Gaussians, the Gaussians are first projected onto the camera plane and then alpha blended with respect to their evaluated density $G(\boldsymbol{x})=\exp\bigl(-\frac{1}{2}(\boldsymbol{x}-\boldsymbol{\mu}_{\mathrm{3D}})^{\top}\boldsymbol{\Sigma}^{-1}(\boldsymbol{x}-\boldsymbol{\mu}_{\mathrm{3D}})\bigr)$ and learned opacity $\alpha$.